OpenStudy (ineedhelp10):

F(x)= -x-4 if x<-7 5x-1 if -7 ≤x≤0 2x+1 if x>0

OpenStudy (ineedhelp10):

@Luigi0210 help please

OpenStudy (freckles):

what do you want to do?

OpenStudy (ineedhelp10):

we're suppose to graph it and find the domain

OpenStudy (freckles):

Ok well let's start with drawing the y=-x-4 for x<-7 part first

OpenStudy (ineedhelp10):

okay and how will that look

OpenStudy (freckles):

that means you are only going to draw y=-x-4 that occurs to the left of x=-7

OpenStudy (freckles):

well we know that y=-x-4 is a line so we only need to know two points on that line before and or at x=-7

OpenStudy (freckles):

if x=-7 then y=-(-7)-4=7-4=3 so plot the point (-7,3) and then plug in some x before x=-7 like x=-8 if x=-8 then y=-(-8)-4=8-4=4 so plot the point (-8,4)

OpenStudy (freckles):

then draw a line connecting those points

OpenStudy (freckles):

but at x=-7 make sure it is a open hole since we have x<-7 and not x=<-7

OpenStudy (freckles):